Peter is 7 years older than his wife ,wendy.The sum.of their ages is 63.Find peter's age
Answer by macston(5194)

.
W=Wendy's age; P=Peter's age=W+7yrs
.
W+P=63yrs
W+W+7yrs=63yrs
2W=56yrs
W=28yrs
Wendy is 28 years old.
.
P=W+7yrs=28yrs+7yrs=35yrs
ANSWER: Peter is 35 years old.
.
CHECK:
W+P=63yrs
28yrs+35yrs=63yrs
63yrs=63yrs
.